At 18C the total volume of a solution formed from MgSO4 and 1.00kg of water fits the expression:
b is understood to be the molaitu of B/b. Calculate the partial molar volumes of the salt and the solvent when in a solution of 0.50 Mol/kg
1. at 18° C the total volume of a solution formed from MgSO4 and 1.00 kg of water fits the expression: V(cm2) 1001.21 +34.96 (b-0.070) where b is understood to be molality in terms of b/bo. Calculate the partial molar volumes of the salt and the solvent when in a solution of 0.050 mol/kg.
Answer #1

partial molar volume of MgSO4 in a mixture is the change in volume per mole of MgSO4 is added to the mixture.

V(cm3)=1001.21+34.96(b-0.070)^2

b=molality=0.050 mol/kg

total volume=V(cm3)= 1001.21+34.96(0.050-0.070)^2=1001.224 cm3

partial molar Volume of water = 1001.21(solvent)

So partial molar volume of MgSO4=1001.224 cm3-1001.21 cm3=0.014 cm3 (salt)